HUBBLE SPACE TELESCOPE DAILY REPORT #2456 PERIOD COVERED: 0000Z (UTC) 09/13/99 - 0000Z (UTC) 09/14/99 Daily Status Report as of 257/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.2 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.3 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed Four Sets of WF/PC-2 8442 (WF/PC-2 Cycle 8 Standard Darks) The WF/PC-2 was used to make a series of Cycle 8 standard dark calibration observations. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD 8218 (Exploring Triton in the Act of Global Change) The Space Telescope Imaging Spectrograph (CCD) was used to make ultraviolet observations of the moon Triton. Triton appears to be undergoing the start of a long-awaited, and perhaps fundamental, global seasonal change. The HST is required for this work owing to Triton's small {0.13"} apparent diameter and the need for ultraviolet observations. The observations were completed as planned, and no problems were reported. 1.6 Completed Three Sets of STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD 8246 (The Youngest Massive Star Clusters in the SMC as Clues to Star Formation in the Early Universe) The Space Telescope Imaging Spectrograph (CCD) was used to make observations the young massive star cluster SMC-N81 that is located in the Small Magellanic Cloud. These are the youngest massive stars accessible to direct observations at ultraviolet and optical wavelengths. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed FGS 8309 (A Link Between Massive Binary Stars and Non-thermal Radio Emission) The Fine Guidance Sensors were used to make astrometric observations of the massive Wolf-Rayet binary star WR125 and several nearby reference stars. Non-thermal radio emission in Wolf-Rayet stars is explained in terms of synchrotron emission from shocks in the stellar wind. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed STIS/CCD 8236 (Black Holes and Gas Disks in a Complete Sample of Radio-Loud Ellipticals - II: Kinematics) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the radio-loud elliptical galaxy NGC-7052. From these observations we will measure the black hole masses, determine the nature, kinematics, and structure of the central gas disks and obtain a modeled upper limit to their mass accretion rates. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.12 Completed WF/PC-2 8274 (A Size Survey of Cometary Nuclei) The WF/PC-2 was to be used to make observations of the comet Forbes as part of a size survey of cometary nuclei. As described in 2.1 and HSTAR 7251, the acquisition for this proposal failed, leaving the take data flag down. 1.13 Completed STIS/CCD 8286 (Absolute Proper Motions of Nearby Dwarf Spheroidal Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the absolute proper motions of the nearby dwarf spheroidal galaxy SCL-QSO-01. HST is essential to achieving the high precision needed to conclusively compare the projected orbital motions of the individual galaxies. Even with our conservative uncertainties, we are competitive with the best ground-based efforts with only a two year baseline. The observations were completed as scheduled, and no anomalies were reported. 1.14 Completed STIS/CCD/MA1 8039 (Spectral Imaging of Stellar Winds) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of star HD4128 (a.k.a., Beta Ceti) in order to make spectral imaging of stellar winds. This is to search for signatures of mass loss in emission lines formed over the wide range of excitation temperatures present in the 1150-1700 A0 interval. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were noted. 1.15 Completed Two Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.16 Completed STIS/CCD/MA1 8241 (A Snapshot Survey of Interstellar Absorption Lines) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make high resolution Echelle spectrographic observations of the standard ultraviolet star HD1909 as part of a snapshot survey of interstellar absorption lines. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed STIS/CCD/MA2 7359 (Chemical Abundances in QSO Broad Absorption Line Regions) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of the quasar Q2116-4439 in order to study the chemical abundances in the QSO’s broad absorption line regions. These observations are designed to test the possible models for high metallicity in BALQSOs and, indirectly, test the origins of the BAL phenomenon. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 6 As described in HSTAR 7251, the FHST updates at 256/122045Z and at 256/122330Z both failed, and the following acquisition also failed. No ARU/PRT was necessary since a roll delay and two full maneuver updates were scheduled prior to the next acquisition. The proposal detailed in 1.12 was affected. The acquisition at 257/070323Z failed when the FGS-3 scan step limit was exceeded. HSTAR 7252 was written. Two proposals, both occurring during the period of the next report (WF/PC 8192 and STIS/CCD 7912), were affected. Scheduled Re-acquisitions: 7 Successful: 7 2.2 FHST Updates: Scheduled: 21 Successful: 19 See comments under 2.1. 2.3 Operations Notes: The SSR EDAC error counter was cleared once per ROP SR-1. The NICMOS FPA temperature sensor for detector 1 was enabled at 256/1816Z. A TTR was written for a non-recoverable 129-second data loss at 257/0136Z, caused by weather conditions at White Sands. ESTR-3 was re-conditioned at 257/0445Z as directed by ROP RD-7. The STIS MCE-1 reset at 257/064936Z while the low voltage was on and while inside an SAA interval. Using ROP NS-12, the STIS flight software error counter was reset at 257/0657Z. MAMA-1 will be recovered at 257/1724Z, the time when the next high voltage will be commanded on via the SMS. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH